OpenStudy (anonymous):

Decompose (3x - 1)/(x^2 - x) into partial fractions

OpenStudy (anonymous):

(3x-1) / (x^2 - x) = (3x-1)/x(x-1) = (A/x) + (B/ x-1). So, A(x-1) + Bx = 3x-1 => Ax-A+Bx= 3x-1 => (A+B)x - A = 3x-1 => A=1. So, A+B =3 => B=2. Finally (3x-1)/ (x^2-x) = [1/x ] + [2/ (x-1)]

OpenStudy (anonymous):

wow umm can u explain??

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

the whole part like idk how to decompose it what is the first step?

OpenStudy (anonymous):

first factor the denominator

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so did you get any roots? for this problem?

OpenStudy (anonymous):

you don't need to find roots for this question

OpenStudy (anonymous):

or solutions either?

OpenStudy (anonymous):

[1/x] + [2/ (x-1)] is the answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

partial fraction means you're breaking down fractions so it's easier to integrate later

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

any more question?

OpenStudy (anonymous):

i'm sorry but that to me looks confusing :/ maybe i need to try it again on my own. thanks tho

OpenStudy (anonymous):

do it step by step. write it down. it's hard at first

OpenStudy (anonymous):

okay i will thanks

OpenStudy (anonymous):

@hcmathkim ok so i did it, can you tell me if its right? 1-3x/x - x^2 = 0 3x - 1/ (x - 1) x = 0 3x - 1/(x - 1) x (multiply both by (x - 1)x 3x - 1 = 0 +1 +1 = 3x = 1 divide 3 both side x = 1/3

OpenStudy (anonymous):

you never said they were equal to zero

OpenStudy (anonymous):

yea its equal to zero...sorry

OpenStudy (anonymous):

then just set the numerator equal to zero. 3x-1=0 => 3x =1 => x=1/3

OpenStudy (anonymous):

okay so what i did up there is correct?

OpenStudy (anonymous):

1) you copied the question wrong 2) you dont need to factor the denominator in this case

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

only way for a fraction to be zero is if the numerator is zero. so just worry about the numerator.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the bottom half of the work is fine

OpenStudy (anonymous):

ok i'll try it again